OpenStudy (anonymous):

Find the slope of the line through the given points. 5. A (1, 5), B (3, 15) D (–2, –4), F (0, –6)

OpenStudy (anonymous):

5

sam (.sam.):

AB= 2 DF=-1

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

A (1, 5), B (3, 15) m = ( 15 -5)/ ( 3 -1) = 10/ 2 = 5 D (–2, –4), F (0, –6) m = ( -6 + 4)/ 2 = -1